OpenStudy (anonymous):

(x/x-2) + (3/x+5) show work please.

OpenStudy (anonymous):

Take the LCM first.. LCM will be : \((x-2)(x+5)\) So: \[\frac{x}{x-2} + \frac{3}{x+5} \implies \frac{x(x+5) + 3(x-2)}{(x-2)(x+5)} \implies \frac{x^2 + 5x + 3x - 6}{(x-2)(x+5)}\] \[\frac{x^2 + 8x - 6}{(x-2)(x+5)}\]

OpenStudy (anonymous):

thanks! i got that too.